To start the construction of a rest house in tagaytay, the engineer advice his foreman to erect five pegs to establish the perimeter (outside) of the floor plan, which is in the shape of a regular pentagon. The desired area of the floor plan is 80 squared meters. Find the length of the string needed in the layout. Consider that 5 percent is accounted for tying.
Answer by Theo(13342)

the formula to use here is taken from the web.

this is what it looks like.



A is the area.
a is the length of a side.

the expression under the square root sign becomes equal to 6.881909602.

the formula becomes A = 1/4 * 6.881909602 * a^2.
solve for a^2 to get a^2 = 4 * A / 6.881909602.
when A = 80, this becomes a^2 = 4 * 80 / 6.881909602 = 46.49872179.
solve for a to get a = square root of 46.49872179 = 6.818997125.
the perimeter of the pentagon is 5 times that = 34.40954801.
add 5% to that to get 35.79973491.
the minimum length of string you'll need has a length of 35.79973491 meters.
that's your solution.

the online calculator that i got this from is found at https://www.google.com/search?q=area+of+a+pentagon+calculator&rlz=1C1CHBD_enUS961US961&oq=area+of+a+pentagon+calculator&gs_lcrp=EgZjaHJvbWUqCQgAEEUYOxiABDIJCAAQRRg7GIAEMgcIARAAGIAEMggIAhAAGBYYHjIICAMQABgWGB4yCAgEEAAYFhgeMggIBRAAGBYYHjIICAYQABgWGB4yCAgHEAAYFhgeMggICBAAGBYYHjIKCAkQABiGAxiKBdIBCDY5MTBqMGo5qAIAsAIA&sourceid=chrome&ie=UTF-8